OpenStudy (anonymous):

Fans & Medals! What is the slope-intercept form of the function that contains the point (–3, 2) and has a slope of –2? y =___x +___

OpenStudy (astrophysics):

Well you need to find the equation of a line which is y = mx+b, where m is the slope and b is where the line crosses the y - axis (y - intercept), so you can use point - slope form and plug in your (x1, y1) points \[ y-y _{1}=m(x-x _{1}) ~~(point~slope~form)\]